OpenStudy (anonymous):

Solve and check 1.8x - 5 = 4.3x - 7.75

OpenStudy (pottersheep):

Rearrange the equation to have all the x's on the same side

OpenStudy (anonymous):

x=1.1

OpenStudy (pottersheep):

Therefore 1.8x - 5 = 4.3x - 7.75 becomes 1.8x - 4.3x = 5 -7.75

OpenStudy (pottersheep):

Remember you switch the +- signs of numbers when you move them accross the equal sign

OpenStudy (pottersheep):

so then you get 6.1x = -2.25 divide both sides by 6.1 to get x

OpenStudy (pottersheep):

x = -.37